OpenStudy (anonymous):

What is the arc length of the subtending arc for an angle of 60° on a circle of radius 3?

OpenStudy (usukidoll):

use this formula \[s=\theta r \] where s = arc theta = degrees r = radius

OpenStudy (usukidoll):

and you also have to convert 60 degrees into radians

OpenStudy (usukidoll):

use \[\frac{\pi}{180} \] to convert from degrees to radians multiply 60 by \[\frac{\pi}{180} \]